7.3 A three-state Markov chain has distinct holding time parameters a, b, and c From each state, the process is equally likely to transition to the other two states. Exhibit the generator matrix and find the stationary distribution.
Answer #1

Here given three state Markov chain with time parameters a, b, c is as follows Matrix

0 1/2 1/2 p=11/2 0 1/2

Let we find stationary distribution is as

0 1/2 1/2 1/2 1/2 0

egin{bmatrix} v_{0} &v_{1} & v_{2} end{bmatrix}= egin{bmatrix} rac{v_{1}}{2}+rac{v_{2}}{2} &rac{v_{0}}{2}+rac{v_{2}}{2} & rac{v_{0}}{2}+rac{v_{1}}{2} end{bmatrix}

that is

v_{0}= rac{v_{1}}{2}+rac{v_{2}}{2} ........................................ (1)

v_{1}= rac{v_{0}}{2}+rac{v_{2}}{2} .........................................(2)

v_{2}=rac{v_{0}}{2}+rac{v_{1}}{2} .......................................... (3)

and V0+ V1 + V2 = 1 ..................................... (a)

From equation (2)

V1 = 0.5(V0 + V2 ) ......................................... (4)

Equation (1) becomes

V0 = 0.5 * (0.5(V0 + V2 )) + 0.5 * V2

V0 = 0.25*V0 +0.25*V2 + 0.5 * V2

0.75*V0 = 0.75 * V2

V0 = V2

Therefore equation (4) becomes

V1 = V0

That is equation (a) becomes

V0 + V0 + V0 = 1

V0 = 1/3

That is Stationary distribution is as

egin{bmatrix} v_{0} &v_{1} & v_{2} end{bmatrix}= egin{bmatrix} rac{1}{3} &rac{1}{3} & rac{1}{3} end{bmatrix}